OpenStudy (anonymous):

how to solve lim x-> sin^2(x/3)/x^2

OpenStudy (anonymous):

First, try direct substitution to see if it will work. This is usually the first step in finding limits.

OpenStudy (anonymous):

how can you solve it for me?

OpenStudy (anonymous):

Did you forget some information when you typed in the limit, because x -> has to have some value after it.

OpenStudy (anonymous):

oh sory x->0

OpenStudy (anonymous):

Okay, then try to replace x=0 into \[\frac{ \sin ^2 (x/3) }{ x^2}\] and see if it gives you a value

OpenStudy (anonymous):

ok w8

OpenStudy (anonymous):

i got sin^2

OpenStudy (anonymous):

In your evaluation, you probably treated sin^2 as a variable. Sin^2 denotes a function (sort of like square root and log), and takes x/3 as its input and outputs a value. Have you taken trig before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Also, as a side note, you also probably canceled x/3=0 and x^2=0. Usually this canceling is fine, but an exception is when both are zero, as in this case.

OpenStudy (anonymous):

Could you please evaluate the expression again, this time keeping in mind that you can't cancel zeroes from the numerator and denominator?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin^{2} (\frac{ x }{ 3 }) / x^2\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \sin^{2} (\frac{ 0 }{ 3 }) / 0^2\]

OpenStudy (anonymous):

I got \[\lim_{x \rightarrow 0} \sin^{2} \] just

OpenStudy (anonymous):

Now what i do? tell me

OpenStudy (zehanz):

If you know this standard limit:\[\lim_{x \rightarrow 0}\frac{ \sin x }{ x } =1\]then it could be that your problem is just the same thing, only slightly changed.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

please reply fast

OpenStudy (zehanz):

When you say it in words: if you calc the quotient of the sine of a number and that number itself, it goes to zero if the number goes to zero. It doesn't have to be x, x/3 will also work:\[\lim_{x \rightarrow 0}\frac{ \sin^2(\frac{ x }{ 3 }) }{ x^2 }=\]\[\left( \lim_{x \rightarrow 0} \frac{ \sin \frac{ x }{ 3 } }{ x }\right)^2\]

OpenStudy (zehanz):

We are almost there: we have the sine of x/3, so wee should divide also by x/3. We only hav x, so if we replace it with x/3, we have to make up for it:\[\left( 3 \cdot \lim_{x \rightarrow 0}\frac{ \sin \frac{ x }{ 3 } }{ \frac{ x }{ 3 } } \right)^2 \]

OpenStudy (zehanz):

See the 3 in front of the limit? Should be 1/3 sorry!

OpenStudy (anonymous):

but the answer is 1/9

OpenStudy (zehanz):

If x goes to 0, so does x/3. If you like you can even replace x/3 by another variable: Suppose u=x/3, then if x -> 0 , also u -> 0 and we get:\[\frac{ 1 }{ 9 }\left( \lim_{u \rightarrow 0} \frac{ \sin u }{ u }\right)^2 \] Now you see why the answer is 1/9: \[\frac{ 1 }{ 9 } \cdot 1^2 =\frac{ 1 }{ 9 }\]

OpenStudy (zehanz):

It is 1/9 because (1/3)^2 =1/9

OpenStudy (anonymous):

how 3 become 1/9

OpenStudy (zehanz):

3 was wrong in the first place. I replaced x with x/3 in the denominator, thereby making the fraction 3 times bigger. To make up for this, I should multiply the fraction with 1/3:\[\left( \frac{ 1 }{ 3 } \lim_{x \rightarrow 0}\frac{ \sin (\frac{ x }{ 3 }) }{ \frac{ x }{ 3 } }\right)^2=\left( \frac{ 1 }{ 3 } \right)^2 \cdot \left( \lim_{u \rightarrow 0} \frac{ \sin u }{ u }\right)^2=\frac{ 1 }{ 9 } \cdot 1^2 =\frac{ 1 }{ 9 }\]I replaced x-3 by u for clarity.

OpenStudy (zehanz):

Aargh! Typo! Last sentence: x/3 = u

OpenStudy (anonymous):

how 1/3 goes their before lim how

OpenStudy (zehanz):

it is a constant:\[\lim_{x \rightarrow a}c \cdot f(x)=c \cdot \lim_{x \rightarrow a} f(x)\]

OpenStudy (anonymous):

ok thanks

OpenStudy (zehanz):

Just remember that in many cases you can get to the standard limit, so alway try to make it look like this:\[\lim_{♣ \rightarrow 0 }\frac{ \sin ♣ }{ ♣ }\]As long as it is three the same things, the outcome is 1.